# Concentration of discrepancy–based approximate Bayesian computation via Rademacher complexity

### v4

@@ 3.2 Concentration as $\varepsilon_{n}\rightarrow\varepsilon^{*}$ and $n\rightarrow\infty$
Theorem 3.2 states our main concentration result. As for related abc theory (see e.g., Proposition 3 in Bernton et al. (2019) and Nguyen et al. (2020)), we also leverage the triangle inequality (2.1). However, we crucially avoid preassuming convergence of $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*})$, and we do not rely on concentration inequalities for $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}}{,}\mu_{\theta})$ regulated by non–explicit sequences of control functions $c_{n}(\theta)f_{n}(\bar{\varepsilon}_{n})$ (Bernton et al., 2019) and $c_{n}(\theta)s_{n}(\bar{\varepsilon}_{n})$ (Nguyen et al., 2020). Rather, we leverage Lemma 2.1 under the proposed Rademacher complexity framework to obtain more direct and informative results, that are also guaranteed to hold uniformly over $\mathcal{P}(\mathcal{Y})$. Indeed, Lemma 2.1 ensures that both $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*})$ and $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\mu_{\theta})$ exceed $2\mathfrak{R}_{n}(\mathfrak{F})$ with a vanishing probability. Therefore, when $\mathfrak{R}_{n}(\mathfrak{F})\rightarrow 0$, we have that $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})\approx\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})$, by combining the triangle inequality in (2.1) with
@@
This means that, if $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})$ is small, then $\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})$ is also small with a high probability. This clarifies the importance of Assumption (IV), which is further supported by the fact that, if $\mathfrak{R}_{n}(\mathfrak{F})$ does not shrink towards zero, by the second inequality in Lemma 2.1, the two empirical measures $\hat{\mu}_{y_{1:n}}$ and $\hat{\mu}_{z_{1:n}}$ do not converge to $\mu^{*}$ and $\mu_{\theta}$, respectively, and hence there is no guarantee that a small discrepancy $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})$ would necessarily imply a vanishing $\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})$.
@@
$\smash{\pi_{n}^{(\varepsilon^{*}+\bar{\varepsilon}_{n})}\Bigl{(}\Bigl{\{}\theta:\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})}>\smash{\varepsilon^{*}+\frac{4\bar{\varepsilon}_{n}}{3}+2\mathfrak{R}_{n}(\mathfrak{F})+\Bigl{[}\frac{2b^{2}}{n}\log\frac{n}{\bar{\varepsilon}_{n}^{L}}\Bigr{]}^{1/2}}\Bigr{\}}\Bigr{)}\leq\frac{2\cdot 3^{L}}{c_{\pi}n},$
@@
The proof of Theorem 3.2 is provided in the Supplementary Material and follows similar arguments considered to establish the abc concentration results in Bernton et al. (2019) and Nguyen et al. (2020), which, in turn, extend those in Frazier et al. (2018). However, as mentioned above, those proofs are specific to a single discrepancy, preassume the convergence of $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*})$, and rely on concentration inequalities for $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\mu_{\theta})$ that depend on non–explicit sequences of control functions. Theorem 3.2 overcomes these issues and proves a unified theory based on the single concentration inequality within Lemma 2.1. This yields a number of technical differences in the proof and, more importantly, it introduces a novel mathematical perspective for the analysis of concentration properties of discrepancy–based abc posteriors. Such a perspective facilitates the derivation of novel theory and proofs (see e.g., Theorem 3.1), along with the extension of current theoretical results to more challenging settings (see Section 6).
@@
Notice that, within Theorem 3.2, the constant $b$ can be typically set equal to $1$ either by definition or upon normalization of the class of $b$–uniformly bounded functions. Moreover, as clarified within the Supplementary Material, Theorem 3.2 also holds when replacing $n/\bar{\varepsilon}_{n}^{L}$ and $c_{\pi}n$ with $M_{n}/\bar{\varepsilon}_{n}^{L}$ and $c_{\pi}M_{n}$, respectively, for any sequence $M_{n}>1$. Nonetheless, to ensure concentration, it suffices to let $M_{n}=n$. In such a case, the quantities $34\bar{\varepsilon}_{n}/$ 3, $2\mathfrak{R}_{n}(\mathfrak{F})$, $[(2b^{2}/n)\log(n/\bar{\varepsilon}_{n}^{L})]^{1/2}$ and $2\cdot 3^{L}/c_{\pi}n$ converge to 00 as $n\rightarrow\infty$, under the settings of Theorem 3.2. This implies that the abc posterior asymptotically concentrates around those $\theta$ values yielding a $\mu_{\theta}$ within discrepancy $\varepsilon^{*}$ from $\mu^{*}$.
@@
In contrast to state–of–the–art theory, this concentration result is guaranteed to hold uniformly over $\mathcal{P}(\mathcal{Y})$, and replaces the currently–employed non–explicit control functions with a known and well–studied quantity, i.e., the Rademacher complexity. Notice that, although Theorem 3.2 provides a unified statement for all ips discrepancies, the bound we derive crucially depends on $\mathfrak{R}_{n}(\mathfrak{F})$, which is specific to each discrepancy $\mathcal{D}_{\mathfrak{F}}$ and plays a fundamental role in controlling the rate of concentration of the abc posterior. In particular, to make the bound as tight as possible, we must choose an $\bar{\varepsilon}_{n}$ such that $34\bar{\varepsilon}_{n}/$ 3 and $[(2b^{2}/n)\log(n/\bar{\varepsilon}_{n}^{L})]^{1/2}$ are of the same order. By neglecting all the terms in $\log\log n$, such a choice leads to setting $\bar{\varepsilon}_{n}$ of the order $[\log(n)/n]^{1/2}$. In this case, the constraint $\bar{\varepsilon}_{n}/\mathfrak{R}_{n}(\mathfrak{F})\to\infty$ may not be satisfied, thereby requiring a larger $\bar{\varepsilon}_{n}$, such as $\mathfrak{R}_{n}(\mathfrak{F})\log\log(n)$. Summarizing, when $\bar{\varepsilon}_{n}=\max\{[\log(n)/n]^{1/2},\mathfrak{R}_{n}(\mathfrak{F})\log\log(n)\}$ it follows, under the conditions of Theorem 3.2, that
$\smash{\pi_{n}^{(\varepsilon^{*}+\bar{\varepsilon}_{n})}(\{\theta:\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})}>\varepsilon^{*}+\mathcal{O}(\bar{\varepsilon}_{n})\})\leq 2\cdot 3^{L}/(c_{\pi}n).$
@@
Notice that the conditions $n\bar{\varepsilon}^{2}_{n}\to\infty$ and $\bar{\varepsilon}_{n}/\mathfrak{R}_{n}(\mathfrak{F})\to\infty$ do not allow to set $\bar{\varepsilon}_{n}<1/\sqrt{n}$. Although this regime is of interest, we are not aware of explicit results in the discrepancy–based abc literature for such a setting. In fact, available studies rely on similar restrictions for some function $f_{n}(\bar{\varepsilon}_{n})$ which is not made explicit except for specific examples that still point toward setting $\bar{\varepsilon}_{n}=[\log(n)/n]^{1/2}>1/\sqrt{n}$ (e.g., Bernton et al., 2019, Supplementary Materials). Faster rates for the abc threshold have been considered by Li and Fearnhead (2018) within the context of summary–based abc, but with a substantially different theoretical focus relative to the one considered here. Recalling the previous discussion, it shall be also emphasized that $\bar{\varepsilon}_{n}<1/\sqrt{n}$ would not yield a faster concentration rate in Theorem 3.2 because of the terms $2\mathfrak{R}_{n}(\mathfrak{F})$ and $[(2b^{2}/n)\log(n/\bar{\varepsilon}_{n}^{L})]^{1/2}$ in the bound.
@@
Theorem 3.2 holds under both well–specified and misspecified models. In the former case, $\mu^{*}=\mu_{\theta^{*}}$ for some $\theta^{*}\in\Theta$. Therefore, $\varepsilon^{*}=0$ and the abc posterior concentrates around those $\theta$ yielding $\mu_{\theta}=\mu^{*}$. Conversely, when the model is misspecified, the abc posterior concentrates on those $\theta$ yielding a $\mu_{\theta}$ within discrepancy $\varepsilon^{*}$ from $\mu^{*}$. Since $\varepsilon^{*}$ is the lowest attainable discrepancy between $\mu_{\theta}$ and $\mu^{*}$, this implies that the abc posterior will concentrate on those $\theta$ such that $\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})=\varepsilon^{*}$. Example A provides an explicit bound for $\varepsilon^{*}$ in a simple yet noteworthy class of misspecified models.
Example A.
@@
(Huber contamination model). In this model, with probability $1-\alpha_{n}$, the data are from a distribution $\mu_{\theta^{*}}$ belonging to the assumed model $\{\mu_{\theta}:\theta\in\Theta\}$, while with probability $\alpha_{n}$ arise from the contaminating distribution $\mu_{C}$. Therefore, the data generating process is $\mu^{*}=(1-\alpha_{n})\mu_{\theta^{*}}+\alpha_{n}\mu_{C},$ with $\alpha_{n}\in[0,1)$ controlling the amount of contamination. In such a context, Definition 2.1 and Assumption (III) imply

### v5

@@ 3.2 Concentration as $\varepsilon_{n}\rightarrow\varepsilon^{*}$ and $n\rightarrow\infty$
Theorem 3.3 states our main concentration result. As in related abc theory (see e.g., Proposition 3 in Bernton et al. (2019) and Nguyen et al. (2020)), we also leverage the triangle inequality (1). However, we crucially avoid pre-assuming convergence of $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*})$, and we do not rely on concentration inequalities for $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}}{,}\mu_{\theta})$ regulated by non-explicit sequences of control functions $c_{n}(\theta)f_{n}(\bar{\varepsilon}_{n})$ (Bernton et al., 2019) and $c_{n}(\theta)s_{n}(\bar{\varepsilon}_{n})$ (Nguyen et al., 2020). Rather, we leverage Lemma 2.6 under the proposed Rademacher complexity framework to obtain more direct and informative results, that are also guaranteed to hold uniformly over $\mathcal{P}(\mathcal{Y})$. Indeed, Lemma 2.6 ensures that both $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*})$ and $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\mu_{\theta})$ exceed $2\mathfrak{R}_{n}(\mathfrak{F})$ with vanishing probability. Therefore, when $\mathfrak{R}_{n}(\mathfrak{F})\rightarrow 0$, we have that $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})\approx\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})$, by combining (1) with
@@
This means that, if $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})$ is small, then $\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})$ is also small with a high probability. This clarifies the importance of Assumption (IV), which is further supported by the fact that, if $\mathfrak{R}_{n}(\mathfrak{F})$ does not shrink to zero, by the second inequality in Lemma 2.6, the two empirical measures $\hat{\mu}_{y_{1:n}}$ and $\hat{\mu}_{z_{1:n}}$ do not converge to $\mu^{*}$ and $\mu_{\theta}$, respectively, and hence there is no guarantee that a small $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\hat{\mu}_{y_{1:n}})$ would imply a vanishing $\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})$.
@@
${\pi_{n}^{(\varepsilon^{*}+\bar{\varepsilon}_{n})}\Bigl{(}\Bigl{\{}\theta:\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})}>{\varepsilon^{*}+\frac{4\bar{\varepsilon}_{n}}{3}+2\mathfrak{R}_{n}(\mathfrak{F})+\Bigl{(}\frac{2b^{2}}{n}\log\frac{n}{\bar{\varepsilon}_{n}^{L}}\Bigr{)}^{1/2}}\Bigr{\}}\Bigr{)}\leq\frac{2\cdot 3^{L}}{c_{\pi}n},$
@@
The proof of Theorem 3.3 can be found in Appendix D of the Supplementary Material and follows similar arguments considered to establish the concentration results in Bernton et al. (2019) and Nguyen et al. (2020), which, in turn, extend those in Frazier et al. (2018). However, as mentioned above, those proofs are specific to a single discrepancy, pre-assume the convergence of $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{y_{1:n}},\mu^{*})$, and rely on concentration inequalities for $\mathcal{D}_{\mathfrak{F}}(\hat{\mu}_{z_{1:n}},\mu_{\theta})$ that depend on non-explicit sequences of control functions. Theorem 3.3 overcomes these issues and proves a unified theory based on the single concentration inequality in Lemma 2.6. This yields technical differences in the proof and, more importantly, it introduces a novel and broadly-impactful perspective for the analysis of concentration properties of discrepancy-based abc posteriors.
@@
Note that in Theorem 3.3 the constant $b$ can be typically set equal to $1$ either by definition or upon normalization of the class of $b$-uniformly bounded functions. Moreover, as clarified in the Supplementary Material, Theorem 3.3 also holds when replacing $n/\bar{\varepsilon}_{n}^{L}$ and $c_{\pi}n$ with $M_{n}/\bar{\varepsilon}_{n}^{L}$ and $c_{\pi}M_{n}$, respectively, for any sequence $M_{n}>1$. Nonetheless, to ensure concentration, it suffices to let $M_{n}=n$. In such a case, the quantities $34\bar{\varepsilon}_{n}/$ 3, $2\mathfrak{R}_{n}(\mathfrak{F})$, $[(2b^{2}/n)\log(n/\bar{\varepsilon}_{n}^{L})]^{1/2}$ and $2\cdot 3^{L}/c_{\pi}n$ converge to 00 as $n\rightarrow\infty$, under the settings of Theorem 3.3. This implies that the abc posterior asymptotically concentrates around those $\theta$ values yielding a $\mu_{\theta}$ within discrepancy $\varepsilon^{*}$ from $\mu^{*}$. Appendix B in the Supplementary Material translates these concentration results from the space of distributions to the space of parameters.
@@
In contrast to available theory, the concentration result in Theorem 3.3 holds uniformly over $\mathcal{P}(\mathcal{Y})$, and replaces the currently-employed non-explicit control functions with a known and well-studied quantity, i.e., the Rademacher complexity. Notice that, although Theorem 3.3 provides a unified statement for all the ips discrepancies, the bound we derive depends on $\mathfrak{R}_{n}(\mathfrak{F})$, which is specific to each discrepancy $\mathcal{D}_{\mathfrak{F}}$ and plays a fundamental role in controlling the rate of concentration of the abc posterior. In particular, to make the bound as tight as possible, we must choose an $\bar{\varepsilon}_{n}$ such that $34\bar{\varepsilon}_{n}/$ 3 and $[(2b^{2}/n)\log(n/\bar{\varepsilon}_{n}^{L})]^{1/2}$ are of the same order. By neglecting all the terms in $\log\log n$, such a choice leads to setting $\bar{\varepsilon}_{n}$ of the order $[\log(n)/n]^{1/2}$. In this case, the constraint $\bar{\varepsilon}_{n}/\mathfrak{R}_{n}(\mathfrak{F})\to\infty$ may not be satisfied, thereby requiring a larger $\bar{\varepsilon}_{n}$, such as $\mathfrak{R}_{n}(\mathfrak{F})\log\log(n)$. Summarizing, when $\bar{\varepsilon}_{n}=\max\{[\log(n)/n]^{1/2},\mathfrak{R}_{n}(\mathfrak{F})\log\log(n)\}$ it follows, under the conditions of Theorem 3.3, that
${\pi_{n}^{(\varepsilon^{*}+\bar{\varepsilon}_{n})}(\{\theta:\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})}>\varepsilon^{*}+\mathcal{O}(\bar{\varepsilon}_{n})\})\leq 2\cdot 3^{L}/(c_{\pi}n).$
@@
Notice that the conditions $n\bar{\varepsilon}^{2}_{n}\to\infty$ and $\bar{\varepsilon}_{n}/\mathfrak{R}_{n}(\mathfrak{F})\to\infty$ do not allow to set $\bar{\varepsilon}_{n}<1/\sqrt{n}$. Although this regime is of interest, we are not aware of explicit results in the discrepancy-based abc literature for such a setting. In fact, available studies rely on similar restrictions for some sequence of functions $f_{n}(\bar{\varepsilon}_{n})$ which is not made explicit except for specific examples that still point toward setting $\bar{\varepsilon}_{n}=[\log(n)/n]^{1/2}>1/\sqrt{n}$ (e.g., Bernton et al., 2019, supplementary materials). Faster rates for the abc threshold have been considered by Li and Fearnhead (2018) in the context of summary-based abc, but with a substantially different theoretical focus and asymptotic regime relative to the one considered here. Recalling, e.g., Theorem 1 in Frazier et al. (2018), within our theory setting the rate $1/\sqrt{n}$ (up to log terms) cannot be improved for summary-based abc; see also Section 4.1 for additional discussion. It shall be also emphasized that $\bar{\varepsilon}_{n}<1/\sqrt{n}$ would not yield a faster concentration rate in Theorem 3.3 because of the terms $2\mathfrak{R}_{n}(\mathfrak{F})$ and $[(2b^{2}/n)\log(n/\bar{\varepsilon}_{n}^{L})]^{1/2}$ in the bound.
@@
Theorem 3.3 holds under both well-specified and misspecified models. In the former case, $\mu^{*}=\mu_{\theta^{*}}$ for some $\theta^{*}\in\Theta$. Therefore, $\varepsilon^{*}=0$ and the abc posterior concentrates around those $\theta$ yielding $\mu_{\theta}=\mu^{*}$. Conversely, when the model is misspecified, the abc posterior concentrates on those $\theta$ yielding a $\mu_{\theta}$ within discrepancy $\varepsilon^{*}$ from $\mu^{*}$. Since $\varepsilon^{*}=\inf\nolimits_{\theta\in\Theta}\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})$, the abc posterior will concentrate on those $\theta$ such that $\mathcal{D}_{\mathfrak{F}}(\mu_{\theta},\mu^{*})$ is arbitrarily close to $\varepsilon^{*}$. Example 3.4 provides an explicit bound for $\varepsilon^{*}$ in a simple yet noteworthy class of misspecified models.
Example 3.4.
@@
(Huber contamination model). In this model, with probability $1-\alpha_{n}$, the data are from a distribution $\mu_{\theta^{*}}$ belonging to the assumed model $\{\mu_{\theta}:\theta\in\Theta\subseteq\mathbb{R}^{p}\}$, while with probability $\alpha_{n}$ arise from a contaminating distribution $\mu_{C}$. Therefore, the data generating process is $\mu^{*}=(1-\alpha_{n})\mu_{\theta^{*}}+\alpha_{n}\mu_{C},$ with $\alpha_{n}\in[0,1)$ controlling the amount of contamination. In such a context, Definition 2.1 and Assumption (III) imply